OpenStudy (anonymous):

I was at the store the other day to get my little brother something for his birthday and so they had a sale on pillow pets for 4/$5 and I WAS trying to figure out how much i will have to pay if i wanted to get 3, and how how will I have to pay to get 1? can someone please show me how you figure this out I totally forgot how too :/!

OpenStudy (angela210793):

4 pillows=5$?

OpenStudy (anonymous):

yes

OpenStudy (angela210793):

so 5/4=1.25$ one pillow

OpenStudy (anonymous):

wait so 5 divided by 4

OpenStudy (angela210793):

yea

OpenStudy (angela210793):

thnx 4 medal :)

OpenStudy (anonymous):

k and how much for 3?

OpenStudy (anonymous):

And np :)

OpenStudy (angela210793):

for 3 y have to multiply 3*1.25=3.75$

OpenStudy (anonymous):

Tysm

OpenStudy (angela210793):

uw:)

OpenStudy (anonymous):

one more thing, can you divide 4*5? or you have to do 5*4?

OpenStudy (angela210793):

u have 4 pillows for 5$ so 4/5 and y'll get the price of 1 pillow

OpenStudy (anonymous):

Ok ty again haha

OpenStudy (angela210793):

:)

OpenStudy (anonymous):

arg sorry again so what would be the price of 3 pillows? >:/

OpenStudy (angela210793):

3*1.25=3.75$ is the price of 3 pillows

OpenStudy (anonymous):

tysm!!!!